If (x-1/2) and (x-22) are the factors of the polynomials px2+5x+r,prove that p=r.
1
Expert's answer
2012-06-07T09:43:29-0400
We get:
px2+5x+r=k(x-1/2)(x-22)=kx2-23k/2+k/11 (k is some, nonzero number)
so p=k and r=k/11 p!=r
Obviously there is a mistake in condition, so the factors got to be (x-1/22) and (x-22), and
then:
px2+5x+r=k(x-1/22)(x-22)=kx2-485k/2+k
so p=k and r=k, and p=r.